# MM-LLMs: Recent Advances in MultiModal Large Language Models

### v4

@@ Abstract
Duzhen Zhang1111Equal contributions. , Yahan Yu2111Equal contributions. , Chenxing Li1 , Jiahua Dong3222Corresponding authors. , Dan Su1, Chenhui Chu2222Corresponding authors. and Dong Yu1 1Tencent AI Lab 2Kyoto University 3Shenyang Institute of Automation, Chinese Academy of Sciences scoutzhang@tencent.com, yahan@nlp.ist.i.kyoto-u.ac.jp
@@ 1 Introduction
With the debut of GPT-4(Vision) OpenAI (2023) and Gemini Team et al. (2023), showcasing impressive MM understanding and generation capabilities, a research fervor on MM-LLMs has been sparked. Initial research primarily focuses on MM content comprehension and text generation, encompassing tasks such as image-text understanding, exemplified by projects like BLIP-2 Li et al. (2023e), LLaVA Liu et al. (2023e), MiniGPT-4 Zhu et al. (2023a), and OpenFlamingo Awadalla et al. (2023); video-text understanding, as demonstrated by initiatives such as VideoChat Li et al. (2023f), Video-ChatGPT Maaz et al. (2023), and LLaMA-VID Li et al. (2023j); and audio-text understanding, as seen in projects like Qwen-Audio Chu et al. (2023b). Later, the capabilities of MM-LLMs have been expanded to support specific modality outputs. This includes tasks with image-text output, such as GILL Koh et al. (2023a), Kosmos-2 Peng et al. (2023), Emu Sun et al. (2024), and MiniGPT-5 Zheng et al. (2023b); as well as speech/audio-text output, exemplified by projects like SpeechGPT Zhang et al. (2023a) and AudioPaLM Rubenstein et al. (2023). Recent research endeavors have focused on mimicking human-like any-to-any modality conversion, shedding light on the path to artificial general intelligence. Some efforts aim to amalgamate LLMs with external tools to reach an approaching any-to-any MM comprehension and generation, such as Visual-ChatGPT Wu et al. (2023a), HuggingGPT Shen et al. (2023), and AudioGPT Huang et al. (2023b). Conversely, to mitigate propagated errors in the cascade system, initiatives like NExT-GPT Wu et al. (2023d), CoDi-2 Tang et al. (2023c), and ModaVerse Wang et al. (2024c) have developed end-to-end MM-LLMs of arbitrary modalities. The timeline of MM-LLMs is depicted in Figure 1.
@@ 2.3 LLM Backbone
Taking LLMs Zhao et al. (2023c); Naveed et al. (2023); Luo et al. (2023) as the core agents, MM-LLMs can inherit some notable properties like zero-shot generalization, few-shot ICL, Chain-of-Thought (CoT), and instruction following. The LLM Backbone processes representations from various modalities, engaging in semantic understanding, reasoning, and decision-making regarding the inputs. It produces (1) direct textual outputs , and (2) signal tokens from other modalities (if any). These signal tokens act as instructions to guide the generator on whether to produce MM contents and, if affirmative, specifying the content to produce:
@@ 3 Training Pipeline
where level=2text width=5em,font=,,
@@
Hong et al. (2023), Osprey Yuan et al. (2023a), V* Wu and Xie (2023), MobileVLM Chu et al. (2023a),
@@
KAM-CoT(Mondal et al.), InternLM-XComposer2 Dong et al. (2024b), MoE-LLaVA Lin et al. (2024a),
@@
MobileVLM V2 Chu et al. (2024), ViGoRYan et al. (2024), VisLingInstruct Zhu et al. (2024b)
@@
3D+TT: 3DMIT Li et al. (2024b)
ManyT: Flamingo Alayrac et al. (2022), MM-REACT Yang et al. (2023b), X-LLM Chen et al. (2023b)
@@
CM3LeonYu et al. (2023a), LaVIT Jin et al. (2024), DreamLLM Dong et al. (2024a), MiniGPT-5
@@
ManyI+T: Emu Sun et al. (2024), BuboGPT() Zhao et al. (2023d), GroundingGPT() Li et al. (2024c)
@@
Gemini Team et al. (2023), ModaVerse Wang et al. (2024c), MLLM-ToolWang et al. (2024a)
@@
(Liu et al.), CogAgent Hong et al. (2023),CLOVA Gao et al. (2023b), -UMi Shen et al. (2024), MLLM-Tool
(Wang et al.), WebVoyager He et al. (2024), Mobile-Agent Wang et al. (2024b)
@@ 3.1 MM PT
The X-Text datasets include Image-Text, Video-Text, and Audio-Text, with Image-Text having two types: Image-Text pairs (e.g., <img1> <txt1>) and interleaved Image-Text corpus (e.g., <txt1><img1><txt2><txt3><img2><txt4>). Details of X-Text datasets are shown in Table 3 of Appendix G.
@@
| DLP | I+TT | I: CLIP/Eva-CLIP ViT | Q-Former+P-Former w/ Linear Projector | OPT/Flan-T5 | – | – | – | – |
@@
| Qwen-VL-(Chat) | I+TT | I: ViT@448 initialized from OpenClip’s ViT-bigG | Cross-attention | Qwen-7B | – | – | 1.4B | 50M |
@@
| MiniGPT-5 | I+TI+T | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-7B | Tiny Transformer w/ MLP | I: StableDiffusion-2 | – | – |
@@
| Lyrics | I+TT | I: CLIP ViT-L/14 & Grounding-DINO-T w/ Swin-T & SAM-HQ w/ MAE & ViT-H & RAM++ w/ Swin-B | MQ-Former w/ Linear Projection | Vicuna-13B | – | – | – | – |
@@ 4 SOTA MM-LLMs
| Qwen-VL | Qwen-7B | – | – | 78.8 | 59.3 | 35.2 | 67.1 | 63.8 | – | – | – | 38.2 | 7.4 | 56.3 | – | – | 59.4 | – | – |
| Qwen-VL-Chat | Qwen-7B | – | – | 78.2 | 57.5 | 38.9 | 68.2 | 61.5 | – | 1487.5 | 360.7 | 60.6 | 56.7 | 58.2 | – | – | – | – | – |
| LLaVA-1.5 | Vicuna-1.5-7B | – | – | 78.5 | 62.0 | 50.0 | 66.8 | 58.2 | 85.9 | 1510.7 | 316.1 | 64.3 | 58.3 | 58.6 | 63.4 | 30.5 | 58.7 | – | – |
@@
| VILA-7B | LLaMA-2-7B | – | – | 79.9 | 62.3 | 57.8 | 68.2 | 64.4 | 85.5 | 1533.0 | – | 68.9 | 61.7 | 61.1 | 69.7 | 34.9 | – | – | – |
| VILA-13B | LLaMA-2-13B | – | – | 80.8 | 63.3 | 60.6 | 73.7 | 66.6 | 84.2 | 1570.1 | – | 70.3 | 64.3 | 62.8 | 73.0 | 38.8 | – | – | – |
| +ShareGPT4V | LLaMA-2-13B | – | – | 80.6 | 63.2 | 62.4 | 73.1 | 65.3 | 84.8 | 1556.5 | – | 70.8 | 65.4 | 61.4 | 78.4 | 45.7 | – | – | – |
@@
5 Benckmarks and Performance
To offer a comprehensive performance comparison, we have compiled a table featuring major MM-LLMs across Vision-Language (VL) benchmarks gathered from various papers Li et al. (2023e); Chen et al. (2023d, f); Lin et al. (2023), shown in Table 2. The information of these benchmarks can be found in Appendix F. Next, we will extract training recipes that boost the effectiveness of MM-LLMs, drawing insights from SOTA models.
@@ 6 Future Directions
More Powerful Models
We can enhance the MM-LLMs’ strength from the following four key avenues: (1) Expanding Modalities: Current MM-LLMs mainly support the following modalities: image, video, audio, 3D, and text. However, the real world involves a broader range of modalities. Extending MM-LLMs to accommodate additional modalities (e.g., web pages, heat maps, and figures&tables) will increase the model’s versatility, making it more universally applicable; (2) Diversifying LLMs: Incorporating various types and sizes of LLMs provides practitioners with the flexibility to select the most appropriate one based on their specific requirements; (3) Improving MM IT Dataset Quality: Current MM IT datasets have ample room for improvement and expansion. Diversifying the range of instructions can enhance the effectiveness of MM-LLMs in understanding and executing user commands; (4) Strengthening MM Generation Capabilities: Most current MM-LLMs are predominantly oriented towards MM understanding. Although some models have incorporated MM generation capabilities, the quality of generated responses may be constrained by the capacities of the LDMs. Exploring the integration of retrieval-based approaches Asai et al. (2023); Gao et al. (2023a) holds significant promise in complementing the generative process, potentially enhancing the overall performance of the model.
@@
Existing benchmarks might not adequately challenge the capabilities of MM-LLMs, given that many datasets have previously appeared to varying degrees in the PT or IT sets. This implies that the models may have learned these tasks during training. Moreover, current benchmarks predominantly concentrate on the VL sub-field. Thus, it is crucial for the development of MM-LLMs to construct a more challenging, larger-scale benchmark that includes more modalities and uses a unified evaluation standard. For instance, GOAT-Bench Lin et al. (2024b) is introduced to assess the capability of various MM-LLMs in discerning and responding to nuanced aspects of social abuse depicted in memes. MathVista Lu et al. (2024) evaluates the math reasoning ability of MM-LLMs within visual contexts. Moreover, MMMU Yue et al. (2023) and CMMMU Zhang et al. (2024) have respectively introduced English and Chinese versions of the massive multi-discipline MM understanding and reasoning benchmark for expert artificial general intelligence. Fan et al. have also challenged MM-LLMs with multipanel VQA. BenchLMM Cai et al. (2023) benchmarks the cross-style visual capability of MM-LLMs. Additionally, Liu et al. have conducted an in-depth study on the optical character recognition capabilities of MM-LLMs.
@@
To deploy MM-LLMs on resource-constrained platforms and achieve optimal performance meanwhile, such as low-power mobile and IoT devices, lightweight implementations are of paramount importance. A notable advancement in this realm is MobileVLM Chu et al. (2023a). This approach strategically downscales LLaMA, allowing for seamless off-the-shelf deployment. MobileVLM further introduces a lightweight downsample projector, consisting of fewer than million parameters, contributing to improved computational speed. Recently, there have been many similar studies on lightweighting MM-LLMs, achieving efficient computation and inference with comparable performance or minimal loss, including TinyGPT-V Yuan et al. (2023b), Vary-toy Wei et al. (2024), Mobile-Agent Wang et al. (2024b), MoE-LLaVA Lin et al. (2024a), and MobileVLM V2 Chu et al. (2024). Nevertheless, this avenue necessitates additional exploration for further advancements in development.
@@
Due to the large training costs associated with their massive scale, MM-LLMs are not amenable to frequent re-training. However, updates are necessary to endow MM-LLMs with new skills and keep them up-to-date with rapidly evolving human knowledge Wu et al. (2024). Thus, Continual Learning (CL) is needed to make the model flexible enough to efficiently and continually leverage emerging data while avoiding the substantial cost of retraining MM-LLMs. CL for MM-LLMs can be classified into two stages: continual PT and continual IT. Recently, a continual MM IT benchmark has been proposed to continuously fine-tune MM-LLMs for new MM tasks while maintaining superior performance on tasks learned during the original MM IT stage He et al. (2023). It introduces two primary challenges: (1) catastrophic forgetting, where models forget previous knowledge when learning new tasks Robins (1995); McCloskey and Cohen (1989); Goodfellow et al. (2013); Zhang et al. (2023d, c, b); Zheng et al. (2023a), and (2) negative forward transfer, indicating that the performance of unseen tasks declines when learning new ones Zheng et al. (2024)

### v5

@@ Abstract
Duzhen Zhang1111Equal contributions.333This work was done when Duzhen Zhang was interning at Tencent, AI Lab, Beijing, China. , Yahan Yu3111Equal contributions. , Jiahua Dong4222Corresponding authors., Chenxing Li1 , Dan Su1, Chenhui Chu3222Corresponding authors. and Dong Yu2 1Tencent AI Lab, China 2Tencent AI Lab, USA 3Kyoto University, Japan 4Mohamed bin Zayed University of Artificial Intelligence, United Arab Emirates {duzhen.zhang972,dongjiahua1995}@gmail.com {yahan@nlp.ist.,chu@}i.kyoto-u.ac.jp, {chenxingli@,dansu@,dyu@global.}tencent.com
@@ 1 Introduction
With the debut of GPT-4(Vision) OpenAI (2023) and Gemini Team et al. (2023), showcasing impressive MM understanding and generation capabilities, a research fervor on MM-LLMs has been sparked. Initial research primarily focuses on MM content comprehension and text generation, encompassing tasks such as image-text understanding, exemplified by projects like BLIP-2 Li et al. (2023e), LLaVA Liu et al. (2023e), MiniGPT-4 Zhu et al. (2023a), and OpenFlamingo Awadalla et al. (2023); video-text understanding, as demonstrated by initiatives such as VideoChat Li et al. (2023f), Video-ChatGPT Maaz et al. (2023), and LLaMA-VID Li et al. (2023j); and audio-text understanding, as seen in projects like Qwen-Audio Chu et al. (2023b). Later, the capabilities of MM-LLMs have been expanded to support specific modality outputs. This includes tasks with image-text output, such as GILL Koh et al. (2023a), Kosmos-2 Peng et al. (2023), Emu Sun et al. (2024), and MiniGPT-5 Zheng et al. (2023b); as well as speech/audio-text output, exemplified by projects like SpeechGPT Zhang et al. (2023a) and AudioPaLM Rubenstein et al. (2023). Recent research endeavors have focused on mimicking human-like any-to-any modality conversion, shedding light on the path to artificial general intelligence. Some efforts aim to amalgamate LLMs with external tools to reach an approaching any-to-any MM comprehension and generation, such as Visual-ChatGPT Wu et al. (2023a), HuggingGPT Shen et al. (2023), and AudioGPT Huang et al. (2023b). Conversely, to mitigate propagated errors in the cascade system, initiatives like NExT-GPT Wu et al. (2023d), CoDi-2 Tang et al. (2023c), and ModaVerse Wang et al. (2024d) have developed end-to-end MM-LLMs of arbitrary modalities. The timeline of MM-LLMs is depicted in Figure 1.
@@ 2.3 LLM Backbone
Taking LLMs Zhao et al. (2023c); Naveed et al. (2023); Mo et al. (2024) as the core agents, MM-LLMs can inherit some notable properties like zero-shot generalization, few-shot ICL, Chain-of-Thought (CoT), and instruction following. The LLM Backbone processes representations from various modalities, engaging in semantic understanding, reasoning, and decision-making regarding the inputs. It produces (1) direct textual outputs , and (2) signal tokens from other modalities (if any). These signal tokens act as instructions to guide the generator on whether to produce MM contents and, if affirmative, specifying the content to produce:
@@ 3 Training Pipeline
where level=2text width=5.15em,font=,,
@@
Hong et al. (2023a), Osprey Yuan et al. (2023a), V* Wu and Xie (2023), MobileVLM Chu et al. (2023a),
@@
KAM-CoT(Mondal et al.), InternLM-XComposer2 Dong et al. (2024c), MoE-LLaVA Lin et al. (2024a),
@@
MobileVLM V2 Chu et al. (2024), ViGoRYan et al. (2024b), VisLingInstruct Zhu et al. (2024b)
@@
3D+TT: 3D-LLM Hong et al. (2023b), Chat-3D Wang et al. (2023e), PointLLM Xu et al. (2023),
3DMIT Li et al. (2024d)
ManyT: Flamingo Alayrac et al. (2022), MM-REACT Yang et al. (2023b), X-LLM Chen et al. (2023b),
@@
CM3LeonYu et al. (2023a), LaVIT Jin et al. (2024b), DreamLLM Dong et al. (2024b), MiniGPT-5
@@
ManyI+T: Emu Sun et al. (2024), BuboGPT() Zhao et al. (2023d), GroundingGPT() Li et al. (2024e)
@@
Gemini Team et al. (2023), ModaVerse Wang et al. (2024d), MLLM-ToolWang et al. (2024b)
@@
(Liu et al.), CogAgent Hong et al. (2023a),CLOVA Gao et al. (2023b), -UMi Shen et al. (2024),MLLM-Tool
(Wang et al.), WebVoyager He et al. (2024), Mobile-Agent Wang et al. (2024c)
@@ 3.1 MM PT
The X-Text datasets include Image-Text, Video-Text, and Audio-Text, with Image-Text having two types: Image-Text pairs (e.g., <img1> <txt1>) and interleaved Image-Text corpus (e.g., <txt1><img1><txt2><txt3><img2><txt4>). Details of X-Text datasets are shown in Table 3.
@@
| DLP | I+TT | I: CLIP/Eva-CLIP ViT | Q-&P-Former w/ Linear Projector | OPT/Flan-T5 | – | – | – | – |
@@
| Qwen-VL-(Chat) | I+TT | I: ViT@448 initialized from OpenClip’s ViT-bigG | Cross-attention | Qwen-7B | – | – | 1.4B† | 50M† |
@@
| MiniGPT-5 | I+TI+T | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-7B | Tiny Transformer | I: StableDiffusion-2 | – | – |
@@
| Lyrics | I+TT | I: CLIP ViT-L/14&Grounding-DINO-T &SAM-HQ&ViT-H&RAM++ | MQ-Former w/ Linear Projection | Vicuna-13B | – | – | – | – |
@@ 4 SOTA MM-LLMs
| Qwen-VL | Qwen-7B | – | – | 78.8∗ | 59.3∗ | 35.2 | 67.1 | 63.8 | – | – | – | 38.2 | 7.4 | 56.3 | – | – | 59.4 | – | – |
| Qwen-VL-Chat | Qwen-7B | – | – | 78.2∗ | 57.5∗ | 38.9 | 68.2 | 61.5 | – | 1487.5 | 360.7 | 60.6 | 56.7 | 58.2 | – | – | – | – | – |
| LLaVA-1.5 | Vicuna-1.5-7B | – | – | 78.5∗ | 62.0∗ | 50.0 | 66.8 | 58.2 | 85.9 | 1510.7 | 316.1‡ | 64.3 | 58.3 | 58.6 | 63.4 | 30.5 | 58.7 | – | – |
@@
| VILA-7B | LLaMA-2-7B | – | – | 79.9∗ | 62.3∗ | 57.8 | 68.2 | 64.4 | 85.5 | 1533.0 | – | 68.9 | 61.7 | 61.1 | 69.7 | 34.9 | – | – | – |
| VILA-13B | LLaMA-2-13B | – | – | 80.8∗ | 63.3∗ | 60.6 | 73.7 | 66.6 | 84.2 | 1570.1 | – | 70.3 | 64.3 | 62.8 | 73.0 | 38.8 | – | – | – |
| +ShareGPT4V | LLaMA-2-13B | – | – | 80.6∗ | 63.2∗ | 62.4 | 73.1 | 65.3 | 84.8 | 1556.5 | – | 70.8 | 65.4 | 61.4 | 78.4 | 45.7 | – | – | – |
@@
5 Benchmarks and Performance
To provide a comprehensive performance comparison, we have compiled a table featuring major MM-LLMs across Vision-Language (VL) benchmarks, as reported in various papers Li et al. (2023e); Chen et al. (2023d, f); Lin et al. (2023). This information is presented in Table 2, with detailed descriptions of these benchmarks available in Appendix F. Given the numerous benchmarks available, we focus on evaluating and comparing different MM-LLMs based on OKVQA, IconVQA, VQA, and GQA.
OKVQA includes questions requiring reasoning with a variety of knowledge types such as commonsense, world knowledge, and visual knowledge. MiniGPT-v2 and MiniGPT-v2-chat perform best in this benchmark, showcasing their outstanding reasoning abilities. IconVQA emphasizes the importance of abstract diagram comprehension and holistic cognitive reasoning in real-world diagram-based word problems, requiring both perceptual acumen and versatile cognitive reasoning. MiniGPT-v2 and MiniGPT-v2-chat also excel in this benchmark, highlighting their exceptional perception and cognitive reasoning capabilities. VQA is a more balanced VQA dataset where each question is paired with a series of images. VILA-13B performs best in this benchmark, demonstrating its superior ability to comprehend multimodal information and its resistance to language biases in the knowledge it acquires. GQA is a VQA dataset focusing on image scene graphs, offering impartial compositional questions derived from real-world images. Each question is associated with a structured representation of its meaning and the detailed logical steps required to answer it. LLaVA-1.5 and VILA-7B perform best in this benchmark, illustrating their excellent reasoning abilities in this domain.
Following this, we will outline training recipes that enhance the effectiveness of MM-LLMs, drawing insights from SOTA models.
@@ 6 Future Directions
More General and Intelligent Models
We can enhance the MM-LLMs’ strength from the following four key avenues: (1) Expanding Modalities: Current MM-LLMs mainly support the following modalities: image, video, audio, 3D, and text. However, the real world involves a broader range of modalities. Extending MM-LLMs to accommodate additional modalities (e.g., web pages, heat maps, and figures&tables) will increase the model’s versatility, making it more universally applicable; (2) Diversifying LLMs: Incorporating various types and sizes of LLMs provides practitioners with the flexibility to select the most appropriate one based on their specific requirements; (3) Improving MM IT Dataset Quality: Current MM IT datasets have ample room for improvement and expansion. Diversifying the range of instructions can enhance the effectiveness of MM-LLMs in understanding and executing user commands; (4) Strengthening MM Generation Capabilities: Most current MM-LLMs are predominantly oriented towards MM understanding. Although some models have incorporated MM generation capabilities, the quality of generated responses may be constrained by the capacities of the LDMs. Exploring the integration of retrieval-based approaches Asai et al. (2023); Gao et al. (2023a); Kang et al. (2024) holds significant promise in complementing the generative process, enhancing the overall performance of the model.
@@
Existing benchmarks may not sufficiently challenge the capabilities of MM-LLMs, as many datasets have appeared to varying degrees in the PT or IT sets. This implies that the models might have already learned these tasks during training. Moreover, current benchmarks predominantly focus on the VL sub-field. Therefore, it is crucial for the development of MM-LLMs to construct a more challenging, larger-scale benchmark that includes additional modalities and employs a unified evaluation standard. For instance, GOAT-Bench Lin et al. (2024b) is designed to assess the capability of various MM-LLMs in discerning and responding to nuanced aspects of social abuse depicted in memes. MMCode Li et al. (2024a) evaluates the algorithmic problem-solving skills of MM-LLMs in visually rich contexts. DecodingTrust Wang et al. (2024a) measures the trustworthiness of MM-LLMs. MathVista Lu et al. (2024) evaluates the mathematical reasoning ability of MM-LLMs within visual contexts, while GeoEval Zhang et al. (2024b); Li et al. (2024f); Song et al. (2024) assesses their proficiency in tackling geometry math problems. Moreover, MMMU Yue et al. (2023) and CMMMU Zhang et al. (2024a) have respectively introduced English and Chinese versions of a comprehensive multi-discipline MM understanding and reasoning benchmark for expert artificial general intelligence. Additionally, Fan et al. have challenged MM-LLMs with multipanel VQA, and BenchLMM Cai et al. (2023) benchmarks the cross-style visual capability of MM-LLMs. Furthermore, Liu et al. have conducted an in-depth study on the optical character recognition capabilities of MM-LLMs. These efforts highlight the need for more sophisticated and diverse benchmarks to truly gauge the advanced capabilities of MM-LLMs.
@@
To deploy MM-LLMs on resource-constrained platforms and achieve optimal performance meanwhile, such as low-power mobile and IoT devices, lightweight implementations are of paramount importance. A notable advancement in this realm is MobileVLM Chu et al. (2023a). This approach strategically downscales LLaMA, allowing for seamless off-the-shelf deployment. MobileVLM further introduces a lightweight downsample projector, consisting of fewer than million parameters, contributing to improved computational speed. Recently, there have been many similar studies on lightweighting MM-LLMs, achieving efficient computation and inference with comparable performance or minimal loss, including TinyGPT-V Yuan et al. (2023b), Vary-toy Wei et al. (2024), Mobile-Agent Wang et al. (2024c), MoE-LLaVA Lin et al. (2024a), and MobileVLM V2 Chu et al. (2024). Nevertheless, this avenue necessitates additional exploration for further advancements in development.
@@
Due to the large training costs associated with their massive scale, MM-LLMs are not amenable to frequent re-training. However, updates are necessary to endow MM-LLMs with new skills and keep them up-to-date with rapidly evolving human knowledge Wu et al. (2024). Thus, Continual Learning (CL) is needed to make the model flexible enough to efficiently and continually leverage emerging data while avoiding the substantial cost of retraining MM-LLMs. CL for MM-LLMs can be classified into two stages: continual PT and continual IT. Recently, a continual MM IT benchmark has been proposed to continuously fine-tune MM-LLMs for new MM tasks while maintaining superior performance on tasks learned during the original MM IT stage He et al. (2023). It introduces two primary challenges: (1) catastrophic forgetting, where models forget previous knowledge when learning new tasks Robins (1995); McCloskey and Cohen (1989); Goodfellow et al. (2013); Zhang et al. (2023d, c, b); Zheng et al. (2023a), and (2) negative forward transfer, indicating that the performance of unseen tasks declines when learning new ones Zheng et al. (2024); Dong et al. (2022, 2024a, 2023b, 2023a).
@@
Biases and Ethical Considerations
Despite the strengths of MM-LLMs, ensuring their safe and efficient application remains crucial. Information generated by MM-LLMs can perpetuate stereotypes and cause harm to vulnerable populations. Since MM-LLMs learn from patterns in MM training data, they can reproduce biases present in these data, potentially leading to representational harm. To address this, we can develop new benchmarks specifically designed to evaluate biases in MM-LLMs Luo et al. (2024). Additionally, designing more effective and fine-grained alignment methods is essential. For instance, using RLHF can help calibrate MM-LLMs to produce answers that align with human values and desires Li et al. (2024c).